# Beyond Fidelity: Explaining Vulnerability Localization of Learning-based Detectors

### 05 Jan 2024 v1

@@ 4. Study Design
In this section, we initially present our evaluation framework concerning the DL-based detectors and the explanation approaches. Subsequently, we elaborate on the evaluation metrics employed for the explanation approaches in subsection 4.2. Lastly, we introduce our pipeline for dataset collection and preprocessing in subsection 4.3.
@@ 4.1. Evaluation Framework
Fig 2 illustrates our evaluation framework for the DL-based detectors and explanation approaches.
@@ 4.1.1. Research Questions (RQ)
We further divide this RQ into two distinct sub-questions, including can explanation approaches perform better on retrained models (RQ3.1)? And whether those detectors learn vulnerability-triggering mechanisms (RQ3.2)?
@@ 4.1.2. Framework Design
In this study, our objective is to explore the vulnerability localization effectiveness of explanation approaches combined with DL-based detectors, as illustrated by Fig 2. We first extract code fragments from collected datasets and label them, perform deduplication and split them into train-set, test-set, and eval-set for further analysis. Then we follow previous work (dwk, ) to employ six widely-used metrics accuracy (ACC), false positive rate (FPR), false negative rate (FNR), recall rate (R), precision (P), and F1 score (F1) to evaluate the performance of the DL-based detectors in identifying vulnerable code fragments. Through initial evaluation, we identify and select the true positive (TP) samples that the detectors successfully detect from test-set.
To evaluate the explanation approaches, we first need to assess whether these explainers can identify important features for the detectors from the samples (RQ1). Subsequently, we evaluate whether the important features selected by the explanation methods correspond to vulnerability-related statements (RQ2). For this purpose, we conduct extensive experiments on TP samples from SARD (Sard, ) and Fan (FanData, ) datasets, obtaining fidelity scores and TLC/FLC scores corresponding to the explanation methods and perform further analysis. We evaluate the explanation methods only on TP samples for two main reasons. First, non-vulnerable code segments are prevalent, rendering vulnerability localization on benign samples not necessary. Second, since the explanation methods operate on top of the detectors’ predictions, vulnerability localization becomes irrelevant if the detector fails to recognize a particular vulnerable code snippet.
@@
Figure 2. Evaluation framework
@@ 4.3.1. Data Source
The dataset used in this study must include annotations for lines of vulnerable code. However, certain real-world datasets like Devign and Reveal only label vulnerable functions (ivdetect, ), which renders them unsuitable for our specific investigation. In the case of the D2A dataset (D2A, ), it labels vulnerable traces in a program using the Infer tool  (Infer, ). Although differential analysis on commit history is employed to reduce false positive rates, many cases still result in false positives. Roland Croft et al. (dataQuality, ) also highlight the presence of over two-thirds of inaccurately labeled samples in D2A, primarily due to false positives in Infer warnings and mislabeled commits. Consequently, we have chosen not to include the D2A dataset in our study.
Therefore, we constructed our dataset by combining two sources: (1) the Software Assurance Reference Dataset (SARD) (Sard, ), a widely utilized vulnerability database, and (2) Fan (FanData, ), which encompasses Common Vulnerabilities and Exposures (CVEs) from 2002 to 2019 and provides 21 features for each vulnerability.
In the SARD dataset, each program or test case can be associated with one or more Common Weakness Enumeration (CWE) IDs, as a program can exhibit multiple vulnerability types. For our study, we follow X Nie et al. (nie2023understanding, ) to primarily focus on six of the top 30 most critical C/C++ software weaknesses in 2021, specifically CWE20, CWE119, CWE125, CWE190, CWE400, and CWE787. To acquire the relevant programs pertaining to these vulnerabilities, we utilize the same crawler employed in DeepWuKong (dwk, ), which enables us to retrieve all the available programs associated with the aforementioned vulnerabilities.
In order to assess the performance of detectors and explainers on complex real-world open-source projects, we conducted additional studies using the Fan dataset. This dataset consists of over 10,000 vulnerable functions along with corresponding fixed code. It is worth noting that Roland Croft et al. (dataQuality, ) have also identified inaccuracies in the labeling of the Fan dataset, primarily attributed to difficulties in accurately tracing the vulnerability-related statements in fixing commits. These inaccuracies could potentially impact the experimental results and should be taken into consideration.
4.3.2. Preprocessing The Benchmark
@@
Labeling SARD. Regarding the labeling of the SARD dataset, the VTS have already been labeled using XML files. After extracting all the code fragments from the original programs, we label each fragment as vulnerable if it contains lines that have been marked as vulnerable in the corresponding XML file, or vice versa. Simultaneously, the vulnerable lines in the vulnerable samples are labeled as triggering lines. However, the VFS in the vulnerable samples of SARD are not labeled. Figure 3 provides an example of a vulnerable code snippet from the SARD dataset, along with its fixed version and labels. It is noteworthy that most vulnerabilities in SARD have a fixed repair mode. Taking this into account, we have implemented an automated labeling mechanism to identify and label the fixing line in the vulnerability code fragments.
Labeling Fan. The Fan dataset is constructed by mining code change information from committed version patches to identify modified lines of code. It includes two types of modifications: added lines and deleted lines. We label the deleted lines in vulnerable functions and the lines in vulnerable functions with data or control dependencies by following the approach used in IVDetect (ivdetect, ). The added lines in the patched version are considered fixed lines corresponding to the vulnerable lines. Notably, triggering lines are not labeled in the Fan dataset.
Figure 3. An example of a vulnerable code fragment and corresponding fixed version with VTS and VFS labelled.
Handling duplicate/conflict samples. Duplicates in the datasets can lead to overfitting and inflated performance metrics. To address this, we follow previous work (dwk, ) to employ a method based on MD5 value comparison to remove duplicate samples. Additionally, there may be samples labeled as both vulnerable and non-vulnerable due to mislabeling in the SARD and Fan datasets. Figure 4 provides an example from the SARD dataset, where a sample is labeled as both vulnerable and non-vulnerable. These samples are typically referred to as benign in a vulnerable context within the “main” functions of SARD when labeled as vulnerable, and in a non-vulnerable context when labeled as non-vulnerable. A similar situation occurs in the Fan dataset, where some code change information is extracted from consecutive patches. Consequently, a fixed code fragment may be modified again in the subsequent patch, resulting in certain functions being labeled as both vulnerable and non-vulnerable. In light of this observation, we label these conflicting samples as vulnerable. In the last, we split each dataset into train-set, eval-set, and test-set with a ratio of 80%-10%-10%.
@@ 5.1. Experimental Setup
For simplicity, we abbreviate DeepWuKong, Reveal, IVDetect, Devign, SySeVR, VulDeePecker, and TokenLSTM as Rev, IVD, Dev, SySe, VDP, and TL in the figures and tables
@@ 5.2. Vulnerability Detector Analysis
In this study, we adopt the provided settings used in the literature or re-implement them for optimal F1 scores. We stop the training process if the F1 score shows no improvements for 10 iterations. The detection results of the DL-base detectors are shown in Fig 5.
Figure 5. Detection performance of seven learning-based detectors in different datasets.
@@ 5.3.1. Fidelity Among Graph-based Explainers
Fig 6 shows the mean experiment results of the graph-based explanation models on four detectors for both SARD and Fan datasets. We observe that:
@@
Fidelity scores of the same explanation method vary significantly across different datasets for a given detector. We can observe that the fidelity+ scores differ between Fan and SARD datasets. This disparity is exemplified in Fig 6 (c.2) and (c.4), which present the results of all five approaches applied to explain IVDetect. Typically, explainers achieve higher fidelity scores by masking more nodes on SARD. However, this trend differs in the Fan dataset which can be illustrated in Fig 6 (b.2) and (b.4) where DeepLift is employed to explain Reveal.
The explainers exhibit varying fidelity scores when interpreting different detectors on identical datasets. For instance, as shown in Fig 6 (b.2) and (d.2), on SARD, The fidelity+ scores of GradCAM on Devign are lower compared to those on Reveal. Similarly, GNNExplainer achieves the highest fidelity+ scores for IVDetect but the lowest for Devign as evidenced in Fig 6 (c.2) and (d.2). This pattern can be observed in other cases as well.
Figure 6. Mean performance of explainers on four detectors over six vulnerability types in SARD and performance on Fan datasets.
Figure 7. Mean performance of explainers on three detectors over six vulnerability types in SARD, and performance on Fan dataset.
@@ 5.3.2. Fidelity Among Sequence-based Explainers
Fig 7 illustrates the fidelity results of five sequence-based explainers for detectors TokenLSTM, SySeVR, and VulDeePecker. Our results bring the following observations:
The fidelity scores of the explainers continue to exhibit variability across different detectors and datasets. Leading to differing comparisons between them. On Sard dataset, GradInput, LRP, Lime respectively reach the highest average fidelity scores on TokenLSTM, SySeVR, VulDeePecker, while Lime, DeepLift respectively has the lowest performance for TokenLSTM, SySeVR and VulDeePecker as shown in Fig 7 (a.1),(b.1),(c.1). On the Fan dataset, Lime emerges as the superior explanation approach among the five with fidelity scores as illustrated in Fig 7 (a.2),(b.2),(c.2).
The explainers generally demonstrate comparatively higher fidelity scores than Random. We can observe the fidelity scores of those explanation approaches in both SARD and Fan are significantly higher than that of Random, with a few exceptional cases. For instance, when explaining SySeVR on SARD, the performance of Random is much lower than the results of the worst-performing explainer DeepLift as shown in Fig 7 (b.1).
@@ 5.4.1. TLC Among Graph-based Explainers
Fig 8 (a) presents the average TLC results for the five graph-based approaches. It allows us to observe the following patterns:
None of these graph-based explainers attain satisfactory TLC scores. We can observe that nearly all methods exhibit a significant deviation from Expectation. While the best results of GradCAM is only about 80% when explaining DeepWuKong as shown in Fig 8 (a.1), it also produces worse results on Reveal as illustrated in Fig 8 (a.2). As for the other four explainers, their performance is worse than Random in most cases.
Analyzing TLC results alongside fidelity scores reveals a weak correlation between the two. For instance, DeepLift achieves the highest fidelity+ scores when explaining DeepWuKong and Reveal as shown in Fig 6 (a.2),(b.2). Yet its TLC is lower than Random for these detectors. Notably, when explaining IVDetect, DeepLift’s TLC scores are significantly lower than Random as illustrated in Fig 8 (a.3). On the other hand, GradCAM performs best in terms of TLC among the five approaches. Still, its fidelity+ scores are notably lower than DeepLift’s when explaining the four detectors as depicted in Fig 6 (a.2),(b.2),(c.2),(d.2).
@@ 5.4.2. FLC Among Graph-based Explainers
Fig 9 displays the experimental results of FLC for the five graph-based explainers on both the SARD and Fan datasets. The following observations can be made:
The FLC results of graph-based explainers on both datasets fall below expectations. On the Fan dataset, the FLC scores of the five explanation approaches on Reveal and Devign are notably lower than Random as shown in Fig 9 (b.2),(b.4), not to mention the comparison with Expectation. When explaining DeepWuKong and IVDetect, DeepLift and GNNExplainer perform the best over other methods respectively. However, their performance still falls considerably short, being even worse than half of Expectation as shown in Fig 9 (b.1),(b.3). On the SARD dataset, DeepLift attains the best FLC scores when explaining DeepWuKong, Reveal, and Devign. Nevertheless, all its top-7 FLC are below 70%. As for GNN-LRP, although it achieves the highest top-7 FLC of 82% when explaining IVDetect as shown in Fig 9 (a.3), there is a significant decrease in performance when explaining the other three detectors.
The results of the explanation methods on FLC show little correlation to TLC and fidelity. On both SARD and Fan dataset, the fidelity+ scores and FLC scores do not align. Despite GNNExplainer achieving the highest fidelity+ scores when explaining IVDetect shown in Fig 6 (c.2), its FLC results are worse than PGExplainer, GradCAM, and GNN-LRP in this case depicted in Fig 9 (a.3). On the Fan dataset, when explaining IVDetect, GNNExplainer performs 1%-4% better than Random as depicted in Fig 9 (b.3). However, there is a significant difference in fidelity+ scores between Random and the other methods as shown in Fig 6 (c.4), indicating a lack of correlation between FLC and fidelity+. Similarly, the FLC scores do not correspond to the TLC scores. While GradCAM achieves the highest TLC scores on DeepWuKong as illustrated in Fig 8 (a.1), its corresponding FLC scores are significantly lower than those of DeepLift and GNN-LRP as shown in Fig 9 (a.1).
@@ 5.4.3. TLC Among Sequence-based Explainers
Fig 8 (b) presents the TLC scores of sequence-based explainers, indicating:
Sequence-based explainers obtain unsatisfactory TLC scores. Similar to the situation in graph-based explainers, the TLC scores of the sequence-based explainer significantly deviate from the Expectation. When explaining SySeVR, Lime produces the highest TLC scores, which are nearly twice than Random as shown in Fig 8 (b.2), however, the scores are even lower than half of the Expectation. The same observation holds true for the other two detectors.
The fidelity scores of sequence-based explainers exhibit a limited correlation with TLC scores. Analyzing the results in terms of fidelity scores, we observe that higher fidelity scores do not necessarily correspond to higher TLC. Taking Lime for example, it attains the highest fidelity scores among five sequence-based explainers as shown in Fig 7 (c.1). However, in regard to TLC scores, it is inferior to DeepLift on VulDeePecker as depicted in Fig 8 (b.3), while DeepLift produces significantly lower fidelity scores than Lime.

### 21 Feb 2024 v2

@@ 4. Study Design
In this section, we first present our evaluation framework concerning the DL-based detectors and the explanation approaches. Subsequently, we elaborate on the evaluation metrics employed for the explanation approaches in § 4.2. Lastly, we introduce our pipeline for dataset collection and preprocessing in § 4.3.
@@ 4.1. Evaluation Framework
Fig. 2 illustrates our evaluation framework for the DL-based detectors and explanation approaches.
@@ 4.1.1. Research Questions (RQ)
We further divide this RQ into two distinct sub-questions, including can explanation approaches perform better on retrained models (RQ3.1)? And can those detectors learn vulnerability-triggering mechanisms (RQ3.2)?
@@ 4.1.2. Framework Design
In this study, our objective is to explore the vulnerability localization effectiveness of explanation approaches combined with DL-based detectors, as illustrated by Fig. 2. We first extract code fragments from collected datasets and label them, perform deduplication and split them into train-set, test-set, and eval-set for further analysis. Then we follow previous work (dwk, ) to employ six widely-used metrics accuracy (ACC), false positive rate (FPR), false negative rate (FNR), recall rate (R), precision (P), and F1 score (F1) to evaluate the performance of the DL-based detectors in identifying vulnerable code fragments. Through initial evaluation, we identify and select the true positive (TP) samples that the detectors successfully detect from test-set.
To evaluate the explanation approaches, we first need to assess whether these explainers can identify important features for the detectors from the samples (RQ1). Subsequently, we evaluate whether the important features selected by the explanation methods correspond to vulnerability-related statements (RQ2). For this purpose, we conduct extensive experiments on TP samples from SARD (Sard, ) and Fan datasets (FanData, ), obtaining fidelity scores and TLC/FLC scores corresponding to the explanation methods and perform further analysis. We evaluate the explanation methods only on TP samples for two main reasons. First, non-vulnerable code segments are prevalent, rendering vulnerability localization on benign samples not necessary. Second, since the explanation methods operate on top of the detectors’ predictions, vulnerability localization becomes irrelevant if the detector fails to recognize a particular vulnerable code snippet.
@@
Figure 2. Evaluation framework.
@@ 4.3.1. Data Source
The dataset used in this study must include annotations for lines of vulnerable code. However, certain real-world datasets like Devign and Reveal only label vulnerable functions (ivdetect, ), which renders them unsuitable for our specific investigation. In the case of the D2A dataset (D2A, ), it labels vulnerable traces in a program using the Infer tool  (Infer, ). Although differential analysis on commit history is employed to reduce false positive rates, many cases still result in false positives. Croft et al. (dataQuality, ) also highlight the presence of over two-thirds of inaccurately labeled samples in D2A, primarily due to false positives in Infer warnings and mislabeled commits. Consequently, we have chosen not to include the D2A dataset in our study.
Therefore, we constructed our dataset by combining two sources: (1) the Software Assurance Reference Dataset (SARD) (Sard, ), a widely utilized vulnerability database, and (2) the Fan dataset (FanData, ), which encompasses Common Vulnerabilities and Exposures (CVEs) from 2002 to 2019 and provides 21 features for each vulnerability.
In the SARD dataset, each program or test case can be associated with one or more Common Weakness Enumeration (CWE) IDs, as a program can exhibit multiple vulnerability types. For our study, we follow Nie et al. (nie2023understanding, ) to primarily focus on six of the top 30 most critical C/C++ software weaknesses in 2021, specifically CWE20, CWE119, CWE125, CWE190, CWE400, and CWE787. To acquire the relevant programs pertaining to these vulnerabilities, we utilize the same crawler employed in DeepWuKong (dwk, ), which enables us to retrieve all the available programs associated with the aforementioned vulnerabilities.
To assess the performance of detectors and explainers on complex real-world open-source projects, we conducted additional studies using the Fan dataset. This dataset consists of over 10,000 vulnerable functions along with corresponding fixed code. It is worth noting that Croft et al. (dataQuality, ) have also identified inaccuracies in the labeling of the Fan dataset, primarily attributed to difficulties in accurately tracing the vulnerability-related statements in fixing commits. These inaccuracies could potentially impact the experimental results and should be taken into consideration.
4.3.2. Preprocessing of the Dataset
@@
Labeling SARD. Regarding the labeling of the SARD dataset, the VTS have already been labeled using XML files. After extracting all the code fragments from the original programs, we label each fragment as vulnerable if it contains lines that have been marked as vulnerable in the corresponding XML file, or vice versa. Simultaneously, the vulnerable lines in the vulnerable samples are labeled as triggering lines. However, the VFS in the vulnerable samples of SARD are not labeled. Fig. 3 provides an example of a vulnerable code snippet from the SARD dataset, along with its fixed version and labels. It is noteworthy that most vulnerabilities in SARD have a fixed repair mode. Taking this into account, we have implemented an automated labeling mechanism to identify and label the fixing line in the vulnerability code fragments.
Labeling Fan. The Fan dataset is constructed by mining code change information from committed version patches to identify modified lines of code. It includes two types of modifications: added lines and deleted lines. We label the deleted lines in vulnerable functions and the lines in vulnerable functions with data or control dependencies by following the approach used in IVDetect (ivdetect, ). The added lines in the patched version are considered fixed lines corresponding to the vulnerable lines. Notably, triggering lines are not labeled in Fan.
Figure 3. Illustration of a vulnerable code fragment and its corresponding fixed version with VTS and VFS labelled.
Handling duplicate/conflict samples. Duplicates in the datasets can lead to overfitting and inflated performance metrics. To address this, we follow previous work (dwk, ) to employ a method based on MD5 value comparison to remove duplicate samples. Additionally, there may be samples labeled as both vulnerable and non-vulnerable due to mislabeling in the SARD and Fan datasets. Fig. 4 provides an example from the SARD dataset, where a sample is labeled as both vulnerable and non-vulnerable. These samples are typically referred to as benign in a vulnerable context within the “main” functions of SARD when labeled as vulnerable, and in a non-vulnerable context when labeled as non-vulnerable. A similar situation occurs in the Fan dataset, where some code change information is extracted from consecutive patches. Consequently, a fixed code fragment may be modified again in the subsequent patch, resulting in certain functions being labeled as both vulnerable and non-vulnerable. In light of this observation, we label these conflicting samples as vulnerable. In the last, we split each dataset into train-set, eval-set, and test-set with a ratio of 80%-10%-10%.
@@ 5.1. Experimental Setup
For simplicity, we abbreviate DeepWuKong, Reveal, IVDetect, Devign, SySeVR, VulDeePecker, and TokenLSTM as DWK, Rev, IVD, Dev, SySe, VDP, and TL in the figures and tables.
@@ 5.2. Vulnerability Detector Analysis
In this study, we adopt the provided settings used in the literature or re-implement them for optimal F1 scores. We stop the training process if the F1 score shows no improvements for 10 iterations. The detection results of the DL-base detectors are shown in Fig. 5.
Figure 5. Detection performance of seven DL-based detectors in different datasets.
@@ 5.3.1. Fidelity Among Graph-based Explainers
Fig. 6 shows the mean experiment results of the graph-based explanation models on four detectors for both SARD and Fan datasets. We observe that:
@@
Fidelity scores of the same explanation method vary significantly across different datasets for a given detector. We can observe that the fidelity+ scores differ between Fan and SARD datasets. This disparity is exemplified in Fig. 6 (c.2) and (c.4), which present the results of all five approaches applied to explain IVDetect. Typically, explainers achieve higher fidelity scores by masking more nodes on SARD. However, this trend differs in the Fan dataset which can be illustrated in Fig. 6 (b.2) and (b.4) where DeepLift is employed to explain Reveal.
The explainers exhibit varying fidelity scores when interpreting different detectors on identical datasets. For instance, as shown in Fig. 6 (b.2) and (d.2), on SARD, The fidelity+ scores of GradCAM on Devign are lower compared to those on Reveal. Similarly, GNNExplainer achieves the highest fidelity+ scores for IVDetect but lower for Devign as evidenced in Fig. 6 (c.2) and (d.2). This pattern can be observed in other cases as well.
Figure 6. Mean fidelity scores of graph-based explainers on four detectors over six vulnerability types in SARD and performance on Fan datasets.
Figure 7. Mean fidelity scores of sequence-based explainers on three detectors over six vulnerability types in SARD, and performance on Fan dataset.
@@ 5.3.2. Fidelity Among Sequence-based Explainers
Fig. 7 illustrates the fidelity results of five sequence-based explainers for detectors TokenLSTM, SySeVR, and VulDeePecker. Our results bring the following observations:
The fidelity scores of the explainers continue to exhibit variability across different detectors and datasets. Leading to differing comparisons between them. On the Sard dataset, GradInput achieves the highest average fidelity scores on TokenLSTM, while Lime attains the highest scores on VulDeePecker and SySeVR. While DeepLift respectively has the lowest performance for TokenLSTM, SySeVR and VulDeePecker as shown in Fig. 7 (a.1),(b.1),(c.1). On the Fan dataset, Lime emerges as the superior explanation approach among the five with fidelity scores as illustrated in Fig. 7 (a.2),(b.2),(c.2).
The explainers generally demonstrate comparatively higher fidelity scores than Random. We can observe the fidelity scores of those explanation approaches in both SARD and Fan are significantly higher than that of Random. For instance, when explaining SySeVR on SARD, the performance of Random is much lower than the results of the worst-performing explainer DeepLift as shown in Fig. 7 (b.1).
@@ 5.4.1. TLC Among Graph-based Explainers
Fig. 8 (a) presents the average TLC results for the five graph-based approaches. It allows us to observe the following patterns:
None of these graph-based explainers attain satisfactory TLC scores. We can observe that nearly all methods exhibit a significant deviation from Expectation. While the best results of GradCAM is only about 80% when explaining DeepWuKong as shown in Fig. 8 (a.1), it also produces worse results on Reveal as illustrated in Fig. 8 (a.2). As for the other four explainers, their performance is worse than Random in most cases.
Analyzing TLC results alongside fidelity scores reveals a weak correlation between the two. For instance, DeepLift achieves the highest fidelity+ scores when explaining DeepWuKong and Reveal as shown in Fig. 6 (a.2),(b.2). Yet its TLC is lower than Random for these detectors. Notably, when explaining IVDetect, DeepLift’s TLC scores are significantly lower than Random as illustrated in Fig. 8 (a.3). On the other hand, GradCAM performs best in terms of TLC among the five approaches. Still, its fidelity+ scores are notably lower than DeepLift’s when explaining the four detectors as depicted in Fig. 6 (a.2),(b.2),(c.2),(d.2).
@@ 5.4.2. FLC Among Graph-based Explainers
Fig. 9 displays the experimental results of FLC for the five graph-based explainers on both the SARD and Fan datasets. The following observations can be made:
The FLC results of graph-based explainers on both datasets fall below expectations. On the Fan dataset, the FLC scores of the five explanation approaches on Reveal and Devign are notably lower than Random as shown in Fig. 9 (b.2),(b.4), not to mention the comparison with Expectation. When explaining DeepWuKong and IVDetect, DeepLift and GNNExplainer perform the best over other methods respectively. However, their performance still falls considerably short, being even worse than half of Expectation as shown in Fig. 9 (b.1),(b.3). On the SARD dataset, DeepLift attains the best FLC scores when explaining DeepWuKong, Reveal, and Devign. Nevertheless, all its top-7 FLC are below 70%. As for GNN-LRP, although it achieves the highest top-7 FLC of 82% when explaining IVDetect as shown in Fig. 9 (a.3), there is a significant decrease in performance when explaining the other three detectors.
The results of the explanation methods on FLC show little correlation to TLC and fidelity. On both SARD and Fan dataset, the fidelity+ scores and FLC scores do not align. Despite GNNExplainer achieving the highest fidelity+ scores when explaining IVDetect shown in Fig. 6 (c.2), its FLC results are worse than PGExplainer, GradCAM, and GNN-LRP in this case depicted in Fig. 9 (a.3). On the Fan dataset, when explaining IVDetect, GNNExplainer performs 1%-4% better than Random as depicted in Fig. 9 (b.3). However, there is a significant difference in fidelity+ scores between Random and the other methods as shown in Fig. 6 (c.4), indicating a lack of correlation between FLC and fidelity+. Similarly, the FLC scores do not correspond to the TLC scores. While GradCAM achieves the highest TLC scores on DeepWuKong as illustrated in Fig. 8 (a.1), its corresponding FLC scores are significantly lower than those of DeepLift and GNN-LRP as shown in Fig. 9 (a.1).
@@ 5.4.3. TLC Among Sequence-based Explainers
Fig. 8 (b) presents the TLC scores of sequence-based explainers, indicating:
Sequence-based explainers obtain unsatisfactory TLC scores. Similar to the situation in graph-based explainers, the TLC scores of the sequence-based explainer significantly deviate from the Expectation. When explaining SySeVR, Lime produces the highest TLC scores, which are nearly twice than Random as shown in Fig. 8 (b.2), however, the scores are even lower than half of the Expectation. The same observation holds true for the other two detectors.
The fidelity scores of sequence-based explainers exhibit a limited correlation with TLC scores. Analyzing the results in terms of fidelity scores, we observe that higher fidelity scores do not necessarily correspond to higher TLC. Taking Lime for example, it attains the highest fidelity scores among five sequence-based explainers as shown in Fig. 7 (c.1). However, in regard to TLC scores, it is inferior to DeepLift on VulDeePecker as depicted in Fig. 8 (b.3), while DeepLift produces significantly lower fidelity scores than Lime.